Circular No. 7658

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001da IN NGC 7780 C. Yu and W. D. Li, University of California at Berkeley, report the dis-

covery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on July 9.5 (mag ∼ 16.5) and 10.5 UT (mag ∼ 16.3). SN 2001da is located at α = 23^{h}53^{m}32^{s}.78, δ = +8^{o}07′02′′.6 (equinox 2000.0), which is 8′′.7 east and 2′′.7 south of the nucleus of NGC 7780. A KAIT image of the same field taken on 2000 Dec. 31.3 showed nothing at this position (limiting mag ∼ 19.0).

μ CENTAURI D. Baade and T. Rivinius, European Southern Observatory (ESO),

Garching; and S. Štefl, Astronomical Institute, Academy of Sciences, Ondbreak řejov, write: “Visual observations made by S. Otero (Liga Iberoamericana de Astronomia, Buenos Aires) and communicated on June 21 imply that the Be star μ Cen (= HR 5193 = HD 120321) brightened between June 7 and June 21 by ∼ 0.2 mag from its previous level of m_{v} = 3.5. An échelle spectrogram taken on June 21.95 UT by X. Liu (University College, London) with the ESO 1.5-m Telescope (+ FEROS) at La Silla shows broad emission wings in Hα and numerous emission lines from singly ionized metals; the peak separations of the Si II line at 634.7 nm and the Fe II lines at II 516.9 and 531.7 nm are 285, 280, and 290 km/s, respectively. Comparison with earlier observations of the same star (Rivinius et al. 1998, A.Ap. 333, 125) suggests that the outburst proper took place ∼ 10 ± 10 days earlier. The ephemeris of Rivinius et al. (1998, ASP Conf. Ser. 135, 343), which is based on the beating of several nonradial pulsation modes, predicts an outburst on June 18 ± 7 days.”

SUPERNOVA 1979C IN NGC 4321 Corrigendum. On IAUC 7657, ‘Supernovae 1993J and 1979C’, last

sentence, for band by the GMRT on 2000 Nov. 9 (36.6 ± 0.6 mJy). read by the GMRT on 2000 Nov. 9 (7.4 ± 0.6 mJy).

SUPERNOVA 2001cz IN NGC 4679 Further to IAUC 7657, R. Chassagne reports that SN 2001cz does not

appear on a CCD image taken on June 25.6 UT (limiting mag ∼ 18).

2001 July 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green